# Gapless symmetry protected topological phases and generalized deconfined critical points from gauging a finite subgroup

### v1

@@ Appendix D Anomaly action
The discussion above about the mixed anomaly between the quotient symmetry and the dual symmetry after gauging a finite subgorup can be generalized to arbitrary dimensions [31]. In $d$-D, the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is $(d-1)$-form, and the anomaly action is given by $\omega=\frac{i}{2}\int_{Y}A^{\hat{\mathbb{Z}}_{2}}\cup dA^{\tilde{U}(1)}$. Here, $A^{\hat{\mathbb{Z}}_{2}}$ is a $d$-form background field. If the symmetry ^$\hat{\mathbb{Z}}_{2}$ again coincides with a UV symmetry $W$ and the gauge theory is emergent as we have discussed in the main text, we then again have a SPT phase with $\alpha=A^{\hat{\mathbb{Z}}_{2}}\cup A^{\mathbb{Z}_{2}}/2,$ which is again canceled by $\omega$. If both $U(1)$ and ^$\hat{\mathbb{Z}}_{2}$ are preserved, then we may have a higher dimensional intrinsically gapless SPT phase. On the other hand, if $\tilde{U}(1)$ is spontaneously broken as in the Higgs phase, the SPT phase coexists with gapless Goldstone modes.

### v2

@@ Appendix D Anomaly action
The discussion above about the mixed anomaly between the quotient symmetry and the dual symmetry after gauging a finite subgroup can be generalized to arbitrary dimensions [31]. In $d$-D, the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is $(d-1)$-form, and the anomaly action is given by $\omega=\frac{i}{2}\int_{Y}A^{\hat{\mathbb{Z}}_{2}}\cup dA^{\tilde{U}(1)}$. Here, $A^{\hat{\mathbb{Z}}_{2}}$ is a $d$-form background field. If the symmetry ^$\hat{\mathbb{Z}}_{2}$ again coincides with a UV symmetry $W$ and the gauge theory is emergent as we have discussed in the main text, we then again have an SPT phase with $\alpha=A^{\hat{\mathbb{Z}}_{2}}\cup A^{\mathbb{Z}_{2}}/2,$ which is again canceled by $\omega$. If both $U(1)$ and ^$\hat{\mathbb{Z}}_{2}$ are preserved, then we may have a higher dimensional intrinsically gapless SPT phase. On the other hand, if $\tilde{U}(1)$ is spontaneously broken as in the Higgs phase, the SPT phase coexists with gapless Goldstone modes.
Appendix E Perturbations by $\sigma^{x}$ and $\sigma^{z}$ in the 1-D Bose-Hubbard model
Here we provide further numerical details on the effects of perturbations in the 1-D Bose-Hubbard model. Specifically, we consider two types of perturbations: $h_{x}\sum_{i}\sigma^{x}_{i+1/2}$, which commutes with the effective gauge constraint from the $K$-term but explicitly breaks the $W$ symmetry, and $h_{z}\sum_{i}\sigma^{z}_{i+1/2}$, which preserves the $W$ symmetry but violates the effective gauge constraint.
Figure 8: Gap scaling with system size when the $\sigma^{x}$ perturbation is added. The $1/L$ behavior already shows up for relatively small system sizes. Parameters used: $t=0.5,U=1,h_{x}=0.1,K=10$.
Figure 9: Gap scaling with system size when the $\sigma^{z}$ perturbation is added. It shows an exponential decay. Parameters used: $t=0.5,U=1,h_{z}=5,K=10$. Here we called $h_{z}$ perturbation, but in fact it has to be of the same order with $K$ for a sizable gap to show up even for small system sizes.
E.1 Perturbation by $\sigma^{x}$
The exact edge degeneracy in the 1-D case discussed in the main text relies crucially on the existence of the $W$ symmetry. Therefore, we would expect the edge degeneracy to disappear once the $W$-breaking $\sigma^{x}$ terms are added to the Hamiltonian. Fig. 8 shows that by adding even a small such perturbation ($h_{x}$ much smaller than the other couplings in the Hamiltonian), the degeneracy is immediately lifted, and the gap follows a $1/L$ scaling, with $L$ being the system size. This means the gapless system no longer have edge modes, hence becomes topologically trivial.
E.2 Perturbation by $\sigma^{z}$
Another interesting type of perturbation is the $\sigma^{z}$ term. While it preserves the $W$ symmetry, it locally anti-commutes with the effective gauge constraint from the large $K$-term. Recall that the localized action of boson parity on the edges, i.e., $P=\sigma^{x}_{1/2}\sigma^{x}_{L+1/2}$, derives from the effective gauge constraint being exactly implemented when $K$ flows to infinity at low energy. Therefore, it is reasonable to expect that if there is a sufficiently large gauge-violating term then $P$ will no longer be exactly localized on the edges, meaning that $P$ can no longer be separated as two independent $\sigma^{x}$ operators in a clean manner. Instead, there will be longer range terms effectively that lead to the hybridization of the two edge modes, splitting the exact degeneracy. This intuition is verified by Fig. 9, which further shows that the gap from degeneracy splitting decays exponentially with system size. Therefore, the effective gauge-violating perturbation could lift the exact degeneracy in the fixed point gapless SPT phase and the original exactly localized edge modes now become exponentially localized.